The age of the father was 4 time and 2 times of his son in 1990 and 2008 respectively. Find the birth year of the son.
Answer by stanbon(75887)

The age of the father was 4 time and 2 times of his son in 1990 and 2008 respectively. Find the birth year of the son.
--------
Equations:
f = 4s
f+18 = 2(s+18)
------
Substitute for "f" and solve for "s"::
4s+18 = 2s+36
2s = 18
s = 9 (son's age now)
----
f = 4s
f = 36 (father's age now)
--------------------------
Cheers,
Stan H.
----------------